### v3

@@ 5.1 Illustration I: Phase Retrieval
We need to specify conditions on the distribution of measurements $(a,b)\sim\mathbb{P}$ that allow us to apply Theorem 5. We recall the concept of a subgaussian tail random vector [44, Section 2.5].
@@
Theorem 6.
@@
$\mathbb{P}^{*}\left(\sup_{x:\left\|{x}\right\|\leq r}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\geq C\sigma r\cdot\max\{\Delta_{\Phi},\Delta_{\Phi}^{2}\}\right)\leq\delta.$ (5.2)
@@
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Notably, for every vector $v$, $\langle a,v\rangle$ is $\left\|{v}\right\|$-subgaussian. Given the property that the product of two subgaussian random variables is subexponential [44, Lemma 2.7.7], we derive that for some universal constant $C>0$, each of
@@
    must be $C\left\|{x_{1}-x_{2}}\right\|$ subexponential for every unit vector $v$ with $\left\|{v}\right\|=1$.
@@
    where $e(x,\xi)=(\psi^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Since $a$ is $\sigma$-subgaussian, $\mathbb{E}[L(\xi)]=2\mathbb{E}[\left\|{a}\right\|^{2}]\leq C\sigma^{2}d$ for some universal constant $C>0$. Thus Assumption C.3 is satisfied with $L(\xi)=2\left\|{a}\right\|^{2}$.
  • (Integrability). Notably $\mathbb{E}[|c(x;\xi)|^{2}]=\mathbb{E}[|b-\langle a,x\rangle|^{2}]<\infty$ for every $x\in\mathbb{R}^{d}$.
@@
=$\mathcal{F}=\left\{\{a\in\mathbb{R}^{d},b\in\mathbb{R}:\langle a,x\rangle^{2}-b\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\right\}.$
We now bound its VC dimension using Theorem 8, whose proof is based on a beautiful theorem in real algebraic geometry [50, 51, 52]. Note $x\mapsto\langle a,x\rangle^{2}-b$ is a degree $2$ polynomial in $x\in\mathbb{R}^{d}$ for every $a,b$. Using Theorem 8, we obtain that for some universal constant $C>0$:
@@
$\mathbb{P}^{*}\left(\sup_{x:x\in\mathbb{B}(0;2r)}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\geq C\sigma r\cdot\max\{\Delta_{\Phi},\Delta_{\Phi}^{2}\}\right)\leq\delta.$ (5.3)
Theorem 6 then follows by recognizing the set inclusion: ${\rm cl}(\mathbb{B}(0;r))\subseteq\mathbb{B}(0;2r)$. ∎
In the literature on optimization, statistics and machine learning, establishing a uniform upper bound on the convergence of subdifferentials is important because it facilitates our understanding of the landscape of the nonsmooth, nonconvex finite sample objective $\phi_{S}$, e.g., the location of stationary points, growth conditions, etc., through the lens of the population objective $\phi$. We shall give further applications in Section 5.1.1 how Theorem 6 contributes to existing results to refine our understanding of the landscape of the nonsmooth phase retrieval objective $\Phi_{S}$ in finite samples.
@@ 5.1.1 Noiseless Phase Retrieval
$\mathcal{Z}_{S}=\{x\in\mathbb{R}^{d}:0\in\partial\Phi_{S}(x)\}~{}~{}~{}\text{and}~{}~{}~{}\mathcal{Z}=\{x\in\mathbb{R}^{d}:0\in\partial\Phi(x)\}.$
@@
In this paper, we show a stronger result for large sample size $m$: the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt{d/m}$ up to logarithmic factors in $d,m$. This rate improvement in $m$ is due to our bound on uniform convergence of subdifferentials of $\Phi$ and $\Phi_{S}$ in Theorem 6, and our proof of Theorem 7 benefits from the characterization of approximate stationary points of the population objective $\Phi$ [14, Corollary 5.3]. Nonetheless, our result, Theorem 7, comes with a tradeoff of additional logarithmic factors in $d$ and $m$.
Theorem 7.
@@
The formal proof of Theorem 7 is deferred to Appendix, Section B.
5.2 Illustration II: Matrix Sensing
Our second illustration demonstrates the flexibility of our framework. Consider the robust matrix sensing objective [10, 11, 41]:
@@ 5.2 Illustration II: Matrix Sensing
Another application of Theorem 5 yields an upper bound on the subdifferential gap for the robust matrix sensing objective. Recall a matrix $A\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian random matrix if the vectorized matrix $\textbf{{\rm vec}}(A)\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian vector in the sense of Definition 3 [44]. For clarity in notation, we use $\left\|{\cdot}\right\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{D\times D}$.
@@
Assume the measurement matrix $A\in\mathbb{R}^{D\times D}$ is $\sigma$-subgaussian, and $\mathbb{E}[|b|^{2}]<\infty$.
@@
The proof is provided in Section C and is analogous to the phase retrieval case. The primary challenge, which comes from evaluating the VC dimension, is addressed using Theorem 8.
6 VC Dimension Bounds
We establish a result on upper bounding the VC dimension of the set $\mathcal{F}$ defined in equation (4.12) when all the functions $x\mapsto c(x;\xi)$ are polynomials. Notably, this result does not require $\xi\mapsto c(x;\xi)$ to be polynomials in the data $\xi$.
Theorem 8.
=$\mathcal{F}=\{\{\xi\in\Xi:c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}.$
is bounded by $Cd\log(Kd)$, where $C>0$ is a universal constant.
Definition 4.
The following theorem, rooted in the field of real algebraic geometry, is referenced from Matousek’s monograph on discrete geometry [52]. See [50, 51] and [53, Theorem 3].
Theorem 9 ([52, Theorem 6.2.1]).
$\left(\frac{50KN}{d}\right)^{d}.$
Proof of Theorem 8.
Consider the set of binary patterns $b$ defined by:
=$\begin{split}\mathcal{B}&=\{(b_{1},b_{2},\ldots,b_{N})\in\{0,1\}^{N}:\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{there exists $z\in\mathbb{R}^{d+1}$ such that}~{}b_{i}=\mathbf{1}\{q_{i}(z)\geq 0\}~{}\text{for all $1\leq i\leq N$}\}.\end{split}$
$|\mathcal{B}|\leq\left(\frac{50KN}{d+1}\right)^{d+1}.$
$\left(\frac{50KN}{d+1}\right)^{d+1}<2^{N}.$
7 Discussions
This paper presents a general technique for proving subdifferential convergence in weakly convex functions through the establishment of subgradient convergence (Theorem 1). We demonstrate this approach to stochastic convex-composite minimizations—an important class of weakly convex objectives—and derive concrete convergence rates for subdifferentials of empirical objectives using tools from statistical learning theories (Theorem 5). Our results achieve optimal rates with tight dependence on the sample size and dimensionality (up to logarithmic factors) and do not rely on key distributional assumptions in the literature requiring the subdifferential mappings to be continuous under the Hausdorff metric. In Section 5, we demonstrate how our results lead to complementary understanding of subdifferential convergence and its uses for landscape analysis, in more specific applications such as robust phase retrieval and matrix sensing problems.
@@ 7 Discussions
  • Can Theorem 1 be extended to other set-valued mappings, such as maximal monotone operators? Extending this theorem could enhance our understanding of solutions to certain stochastic variational inequalities [56].
  • Theorem 5 could facilitate our understanding of the landscape of a certain class of nonsmooth empirical risk minimizations. Exploring its statistical applications, such as nonlinear quantile regressions [13, 57], may lead to new insights into data-driven models.
@@
Acknowledgements Feng Ruan extends his gratitude to his friends and advisor—John Duchi, Qiyang Han, X.Y. Han, Kate Yao and Yichen Zhang—for their support and patience as he tackled this once seemingly challenging problem. A shift in perspective during his time at Northwestern led him to this solution. He is also grateful to Damek Davis and Dmitriy Drusvyatskiy for their valuable feedback on the initial manuscript draft posted on arXiv.
@@ Appendix A Proof of Theorem 4
In this section, we prove Theorem 4. The key to our proof is to establish high probability upper bounds of the following two quantities:
$\begin{split}\Gamma_{1}&=\sup_{x\in\mathbb{B}(x_{0};r),t\in\mathbb{R}}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x;\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x;\xi)]}\right\|\\ \Gamma_{2}&=\sup_{x\in\mathbb{B}(x_{0};r)}~{}~{}~{}~{}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[(h^{sm})^{\prime}(c(x;\xi))\cdot\nabla c(x;\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[(h^{sm})^{\prime}(c(x;\xi))\cdot\nabla c(x;\xi)]}\right\|\end{split}$
Our interests on these two quantities $\Gamma_{1},\Gamma_{2}$ are motivated by the following inequality:
@@
Below, we will establish the existence of a universal constant $C>0$ such that for every $\delta\in(0,1)$:
$\mathbb{P}^{*}(\Gamma_{1}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta,~{}~{}~{}\mathbb{P}^{*}(\Gamma_{2}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta$ (A.2)
where $\Delta$ has the same definition as in equation (4.16):
Given these high probability bounds on $\Gamma_{1},\Gamma_{2}$ in equation (A.2), Theorem 4 follows by plugging these high probability bounds into equation (A.1).
It remains to prove the high probability bounds on $\Gamma_{1}$ and $\Gamma_{2}$ as specified in equation (A.2). To unify the proofs for these two bounds, we utilize Theorem 10, which details the uniform convergence rate involving indicator functions and smooth functions. The proof of Theorem 10 is in Section A.1.
@@
Consider two subsets of $\mathbb{R}^{d}$, $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, where $\mathcal{Z}_{2}$ is a Euclidean ball centered at $z_{2,0}$ with radius $r$. Let $o:\mathcal{Z}_{1}\times\Xi\to\mathbb{R}$ and $w:\mathcal{Z}_{2}\times\Xi\to\mathbb{R}^{d}$. Suppose $\xi_{1},\xi_{2},...,\xi_{m}$ are i.i.d. samples from a distribution $\mathbb{P}$ with support $\Xi$. For any $z_{1}\in\mathcal{Z}_{1}$ and $z_{2}\in\mathcal{Z}_{2}$, we define:
$\begin{split}K_{m}(z_{1},z_{2},t)&=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}\left\{o(z_{1},\xi_{i})\geq t\right\}w(z_{2},\xi_{i}),~{}~{}~{}~{}\\ K(z_{1},z_{2},t)&=\mathbb{E}_{\xi\sim\mathbb{P}}\bigg{[}\mathbf{1}\left\{o(z_{1},\xi)\geq t\right\}w(z_{2},\xi)\bigg{]}\end{split}.$
@@
    The random function $z_{2}\mapsto w(z_{2},\xi)$ and $z_{1}\mapsto o(z_{1},\xi)$ are continuous.
@@
Here, $\mathcal{W}:=\left\{\{\xi\in\Xi:o(z_{1},\xi)\geq t\}\mid z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}\right\}$, and ${\rm vc}(\mathcal{W})$ denotes its VC dimension.
We will apply Theorem 10 for two different sets of functions $o,w$ to deduce the high probability bounds on $\Gamma_{1},\Gamma_{2}$ in equation (A.2). Below we set $\mathcal{Z}_{1}=\mathbb{R}^{d}$ and $\mathcal{Z}_{2}=\mathbb{B}(x_{0};r)$.
First, we set $o(x,\xi)=c(x,\xi)$ and $w(x^{\prime},\xi)=\nabla c(x^{\prime},\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem 10 are satisfied. Applying this theorem yields the following guarantee. Let
$\begin{split}\Gamma_{1}^{\prime}=\sup_{x\in\mathbb{R}^{d},t\in\mathbb{R},x^{\prime}\in\mathbb{B}(x_{0};r)}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x^{\prime};\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x^{\prime};\xi)]}\right\|\end{split}$
@@
$\mathbb{P}^{*}(\Gamma_{1}^{\prime}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta,$
where $\Delta$ is defined as in equation (4.16). Since $\Gamma_{1}\leq\Gamma_{1}^{\prime}$ by definition, this establishes the first inequality of equation (A.2).
Next, we set $o(x,\xi)\equiv 0$ and $w(x^{\prime},\xi)=(h^{sm})^{\prime}(c(x^{\prime};\xi))\cdot\nabla c(x^{\prime};\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem 10 are satisfied. Applying this theorem provides the following guarantee. Let
$\begin{split}\Gamma_{2}^{\prime}=\sup_{t\in\mathbb{R},x^{\prime}\in\mathbb{B}(x_{0};r)}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\mathbf{1}\{0\geq t\}\cdot(h^{sm})^{\prime}(c(x^{\prime};\xi))\nabla c(x^{\prime};\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[\mathbf{1}\{0\geq t\}\cdot(h^{sm})^{\prime}(c(x^{\prime};\xi))\nabla c(x^{\prime};\xi)]}\right\|\end{split}$
$\mathbb{P}^{*}(\Gamma_{2}^{\prime}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta,$
where $\Delta$ is defined as in equation (4.16). Since $\Gamma_{2}\leq\Gamma_{2}^{\prime}$ by definition, this establishes the second inequality of equation (A.2).
At this point, the proof of Theorem 4 is complete.

### v4

@@ 6.1 Illustration I: Phase Retrieval
We need to specify conditions on the distribution of measurements $(a,b)\sim\mathbb{P}$ that allow us to apply Theorem 5. We recall the concept of a subgaussian tail random vector [46, Section 2.5].
@@
Corollary 1.
Assume the measurement vector $a$ is $\sigma$-subgaussian, and $\mathbb{E}[|b|]<\infty$.
@@
$\mathbb{P}^{*}\left(\sup_{x:\left\|{x}\right\|\leq r}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\geq C\sigma^{2}r\cdot\max\{\Delta_{\Phi},\Delta_{\Phi}^{2}\}\right)\leq\delta.$ (6.2)
@@
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Notably, for every vector $v$, $\langle a,v\rangle$ is $\sigma\left\|{v}\right\|$-subgaussian. Given the property that the product of two subgaussian random variables is subexponential [46, Lemma 2.7.7], we derive that for some universal constant $C>0$, each of
@@
    must be $C\sigma^{2}\left\|{x_{1}-x_{2}}\right\|$ subexponential for every unit vector $v$ with $\left\|{v}\right\|=1$.
@@
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Since $a$ is $\sigma$-subgaussian, $\mathbb{E}[L(\xi)]=2\mathbb{E}[\left\|{a}\right\|^{2}]\leq C\sigma^{2}d$ for some universal constant $C>0$. Thus Assumption C.3 is satisfied with $L(\xi)=2\left\|{a}\right\|^{2}$.
  • (Integrability). Notably $\mathbb{E}[h(c(x;\xi))]=\mathbb{E}[|b-\langle a,x\rangle^{2}|]<\infty$ for every $x\in\mathbb{R}^{d}$.
@@
=$\mathcal{F}=\left\{\{a\in\mathbb{R}^{d},b\in\mathbb{R}\colon\langle a,x\rangle^{2}-b\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\right\}.$
We now bound its VC dimension using Theorem 6. Note $x\mapsto\langle a,x\rangle^{2}-b$ is a degree $2$ polynomial in $x\in\mathbb{R}^{d}$ for every $a,b$. Using Theorem 6, we obtain that for some universal constant $C>0$:
@@
$\mathbb{P}^{*}\left(\sup_{x:x\in\mathbb{B}(0;2r)}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\geq C\sigma^{2}r\cdot\max\{\Delta_{\Phi},\Delta_{\Phi}^{2}\}\right)\leq\delta.$ (6.3)
Corollary 1 then follows by recognizing the set inclusion: ${\rm cl}(\mathbb{B}(0;r))\subseteq\mathbb{B}(0;2r)$. ∎
In the literature on optimization, statistics and machine learning, establishing a uniform upper bound on the convergence of subdifferentials is important because it facilitates our understanding of the landscape of the nonsmooth, nonconvex finite sample objective $\phi_{S}$, e.g., the location of stationary points, growth conditions, etc., through the lens of the population objective $\phi$. We shall give further applications in Section 6.1.1 how Corollary 1 contributes to existing results to refine our understanding of the landscape of the nonsmooth phase retrieval objective $\Phi_{S}$ in finite samples.
@@ 6.1.1 Noiseless Phase Retrieval
$\mathcal{Z}_{S}=\{x\in\mathbb{R}^{d}\colon 0\in\partial\Phi_{S}(x)\}~{}~{}~{}\text{and}~{}~{}~{}\mathcal{Z}=\{x\in\mathbb{R}^{d}\colon 0\in\partial\Phi(x)\}.$
@@
In this paper, we show a stronger result for large sample size $m$: the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt{d/m}$ up to logarithmic factors in $d,m$. This rate improvement in $m$ is due to our bound on uniform convergence of subdifferentials of $\Phi$ and $\Phi_{S}$ in Corollary 1, and our proof of Proposition 8 benefits from the characterization of approximate stationary points of the population objective $\Phi$ [14, Corollary 5.3]. Nonetheless, our result, Proposition 8, comes with a tradeoff of additional logarithmic factors in $d$ and $m$.
Proposition 8.
@@
The formal proof of Proposition 8 is deferred to Appendix, Section B.
6.2 Illustration II: Blind Deconvolution
Our second example studies the (real-valued) blind deconvolution objective:
$\begin{split}\min_{y,w}\tilde{\Phi}(y,w)~{}~{}&\text{where}~{}~{}\tilde{\Phi}(y,w)=\mathbb{E}_{(u,v,b)\sim\mathbb{P}}|\langle u,y\rangle\langle v,w\rangle-b|.\\ \min_{y,w}\tilde{\Phi}_{S}(y,w)~{}~{}&\text{where}~{}~{}\tilde{\Phi}(y,w)=\mathbb{E}_{(u,v,b)\sim\mathbb{P}_{m}}|\langle u,y\rangle\langle v,w\rangle-b|=\frac{1}{m}\sum_{i=1}^{m}|\langle u_{i},y\rangle\langle v_{i},w\rangle-b_{i}|.\end{split}$
This objective proposed in the literature [9, 43] falls into the stochastic convex-composite minimization problems in Section 4, as we can simply set $x=(y,w)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}$, $\xi=(u,v,b)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\times\mathbb{R}$, $h(z)=|z|$ and $c(x;\xi)=\langle u,y\rangle\langle v,w\rangle-b$.
Another application of Theorem 5 yields an upper bound on the subdifferential gap for the robust blind deconvolution objective.
Corollary 2.
Assume the measurements $u,v$ are independent $\sigma$-subgaussian vectors, and $\mathbb{E}[|b|]<\infty$.
$\mathbb{P}^{*}\left(\sup_{x:\left\|{x}\right\|\leq r}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\geq C\sigma^{2}r\cdot\max\{\Delta_{\tilde{\Phi}},\Delta_{\tilde{\Phi}}^{2}\}\right)\leq\delta.$ (6.4)
In the above,
$\Delta_{\tilde{\Phi}}=\sqrt{\frac{1}{m}\cdot\left((d_{1}+d_{2})\log(d_{1}+d_{2})\log m+\log(\frac{1}{\delta})\right)}.$
We wish to comment that, the recent work [43] has studied the location of the stationary points of the finite sample objective $\tilde{\Phi}_{S}$ for the noiseless blind deconolution problem, assuming independent measurements $u\sim\mathsf{N}(0,I_{d_{1}})$, and $v\sim\mathsf{N}(0,I_{d_{2}})$, and no corruptions in the measurement outcome $b=\langle u,\bar{y}\rangle\langle v,\bar{w}\rangle$. Their analysis employs a version of Attouch’s epi-convergence theorem, similar to its use in the noiseless phase retrieval problem [14], to identify the location of finite sample stationary points. In Section 6.1.1, we demonstrated that our approach achieves tighter convergence rates in the context of phase retrieval. Although we expect that similar improvements can be achieved for blind deconvolution, a full analysis remains to be explored and is beyond the scope of the current paper.
6.3 Illustration III: Matrix Sensing
Our final example demonstrates the flexibility of our framework. Consider the robust matrix sensing objective [10, 11, 42]:
@@ 6.3 Illustration III: Matrix Sensing
Another application of Theorem 5 yields an upper bound on the subdifferential gap for the robust matrix sensing objective. Recall a matrix $A\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian random matrix if the vectorized matrix $\textbf{{\rm vec}}(A)\in\mathbb{R}^{D\times D}$ is a $\sigma$-subgaussian vector in the sense of Definition 4 [46]. Below, we use $\left\|{\cdot}\right\|_{F}$ to denote the Frobenius norm of a matrix in $\mathbb{R}^{D\times D}$.
@@
Assume the measurement matrix $A\in\mathbb{R}^{D\times D}$ is $\sigma$-subgaussian, and $\mathbb{E}[|b|]<\infty$.
@@
The proof is provided in Section D and is analogous to the phase retrieval case. The primary challenge, which comes from evaluating the VC dimension, is addressed using Theorem 6.
7 Discussion
This paper presents a general technique for proving subdifferential convergence in weakly convex functions through the establishment of subgradient convergence (Theorem 1). We demonstrate this approach to stochastic convex-composite minimization problems—an important class of weakly convex objectives—and derive concrete convergence rates for subdifferentials of empirical objectives using tools from statistical learning theories (Theorem 5). Our results achieve optimal rates with tight dependence on the sample size and dimensionality (up to logarithmic factors) and do not rely on key distributional assumptions in the literature requiring the population objectives to be continuously differentiable. In Section 6, we demonstrate how our results lead to complementary understanding of subdifferential convergence and its uses for landscape analysis, in more specific applications such as robust phase retrieval and matrix sensing problems.
@@ 7 Discussion
  • Can Theorem 1 be utilized to deduce concrete convergence rates for other interesting class of stochastic locally Lipschitz minimization problems? A perhaps interesting example is the convergence of subdifferentials in stochastic difference-of-convex optimization problems [58].
  • Theorem 5 could facilitate our understanding of the landscape of a certain class of nonsmooth empirical risk minimization problems. Exploring its statistical applications, such as nonlinear quantile regressions [13, 59], may lead to new insights into data-driven models.
@@
Acknowledgements Feng Ruan extends his gratitude to his friends and advisor—John Duchi, Qiyang Han, X.Y. Han, Kate Yao and Yichen Zhang—for their support and patience as he tackled this once seemingly challenging problem. A shift in perspective during his time at Northwestern led him to this solution. He is also grateful to Damek Davis and Dmitriy Drusvyatskiy for their valuable feedback on the initial manuscript draft posted on arXiv. Feng Ruan would also like to express gratitude to Alexander Shapiro, and the anonymous reviewers for their valuable discussions on literature and insights regarding the proof of Theorem 1 using variational analytic tools.
@@ Appendix A Proof of Theorem 4
Our proof starts with a basic triangle inequality bound:
@@
where the two quantities $\Gamma_{1},\Gamma_{2}$ are defined by:
$\begin{split}\Gamma_{1}&=\sup_{x\in\mathbb{B}(x_{0};r),t\in\mathbb{R}}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x;\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x;\xi)]}\right\|\\ \Gamma_{2}&=\sup_{x\in\mathbb{B}(x_{0};r)}~{}~{}~{}~{}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[(h^{sm})^{\prime}(c(x;\xi))\cdot\nabla c(x;\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[(h^{sm})^{\prime}(c(x;\xi))\cdot\nabla c(x;\xi)]}\right\|\end{split}.$
We will deduce high probability upper bounds on $\Gamma_{1}$ and $\Gamma_{2}$ using Theorem 9. The proof of Theorem 9 is technical and lengthy, and deferred to Section A.1.
@@
Consider two subsets of $\mathbb{R}^{d}$, $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$, where $\mathcal{Z}_{2}$ is a Euclidean ball centered at $z_{2,0}$ with radius $r$. Let $q\colon\mathcal{Z}_{1}\times\Xi\to\mathbb{R}$ and $w\colon\mathcal{Z}_{2}\times\Xi\to\mathbb{R}^{d}$. Suppose $\xi_{1},\xi_{2},...,\xi_{m}$ are i.i.d. samples from a distribution $\mathbb{P}$ with support $\Xi$. For any $z_{1}\in\mathcal{Z}_{1}$ and $z_{2}\in\mathcal{Z}_{2}$, we define:
$\begin{split}K_{m}(z_{1},z_{2},t)&=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}\left\{q(z_{1},\xi_{i})\geq t\right\}w(z_{2},\xi_{i}),~{}~{}~{}~{}\\ K(z_{1},z_{2},t)&=\mathbb{E}_{\xi\sim\mathbb{P}}\bigg{[}\mathbf{1}\left\{q(z_{1},\xi)\geq t\right\}w(z_{2},\xi)\bigg{]}\end{split}.$
@@
    The random functions $z_{2}\mapsto w(z_{2},\xi)$ and $z_{1}\mapsto q(z_{1},\xi)$ are continuous.
@@
Here, $\mathcal{W}:=\left\{\{\xi\in\Xi:q(z_{1},\xi)\geq t\}\mid z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}\right\}$, and ${\rm vc}(\mathcal{W})$ denotes its VC dimension.
We go back and prove Theorem 4.
We deduce high probability bounds on $\Gamma_{1},\Gamma_{2}$ by applying Theorem 9 to two different sets of functions $q,w$. Recall the following definition of $\Delta$ as given in equation (4.16):
Below we set $\mathcal{Z}_{1}=\mathbb{R}^{d}$ and $\mathcal{Z}_{2}=\mathbb{B}(x_{0};r)$.
  1. 1.
    (Bounds on $\Gamma_{1}$) First, we set $q(x,\xi)=c(x,\xi)$ and $w(x^{\prime},\xi)=\nabla c(x^{\prime},\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem 9 are satisfied. Applying this theorem yields the following guarantee. Let us define
    $\begin{split}\Gamma_{1}^{\prime}=\sup_{x\in\mathbb{R}^{d},t\in\mathbb{R},x^{\prime}\in\mathbb{B}(x_{0};r)}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x^{\prime};\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[\mathbf{1}\{c(x;\xi)\geq t\}\cdot\nabla c(x^{\prime};\xi)]}\right\|.\end{split}$
@@
    $\mathbb{P}^{*}(\Gamma_{1}^{\prime}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta.$
    Since $\Gamma_{1}\leq\Gamma_{1}^{\prime}$ always holds by definition, this establishes the bound:
    $\mathbb{P}^{*}(\Gamma_{1}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta.$ (A.3)
  2. 2.
    (Bounds on $\Gamma_{2}$) Next, we set $q(x,\xi)\equiv 0$ and $w(x^{\prime},\xi)=(h^{sm})^{\prime}(c(x^{\prime};\xi))\cdot\nabla c(x^{\prime};\xi)$. Given Assumptions C.1 and C.2, the conditions of Theorem 9 are satisfied. Applying this theorem provides the following guarantee. Let us define
    $\begin{split}\Gamma_{2}^{\prime}=\sup_{t\in\mathbb{R},x^{\prime}\in\mathbb{B}(x_{0};r)}\left\|{\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\mathbf{1}\{0\geq t\}\cdot(h^{sm})^{\prime}(c(x^{\prime};\xi))\nabla c(x^{\prime};\xi)]-\mathbb{E}_{\xi\sim\mathbb{P}}[\mathbf{1}\{0\geq t\}\cdot(h^{sm})^{\prime}(c(x^{\prime};\xi))\nabla c(x^{\prime};\xi)]}\right\|.\end{split}$
    $\mathbb{P}^{*}(\Gamma_{2}^{\prime}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta.$
    Since $\Gamma_{2}\leq\Gamma_{2}^{\prime}$ by definition, this establishes the bound:
    $\mathbb{P}^{*}(\Gamma_{2}\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\})\leq\delta.$ (A.4)
With both high probability bounds on $\Gamma_{1}$ and $\Gamma_{2}$ in equations (A.3) and (A.4), we can use the inequality (A.1) to deduce that
$\mathbb{P}^{*}\left(\sup_{x\in\mathbb{B}(x_{0};r)}\left\|{G_{S}(x)-G(x)}\right\|\geq C\cdot\Big{(}\sum_{j=1}^{\infty}a_{j}+1\Big{)}\cdot(\sigma r+\sigma_{0})\cdot\max\{\Delta,\Delta^{2}\}\right)\leq 2\delta.$
By appropriately adjusting the numerical constant from $\delta$ to $\delta/2$, and noting that $\zeta=1+\sum_{j=1}^{\infty}a_{j}$ by definition, Theorem 4 follows.